# Controllable Latent Diffusion for Traffic Simulation

### v1

@@ Abstract
The validation of autonomous driving systems benefits greatly from the ability to generate scenarios that are both realistic and precisely controllable. Conventional approaches, such as real-world test drives, are not only expensive but also lack the flexibility to capture targeted edge cases for thorough evaluation. To address these challenges, we propose a controllable latent diffusion that guides the training of diffusion models via reinforcement learning to automatically generate a diverse and controllable set of driving scenarios for virtual testing. Our approach removes the reliance on large-scale real-world data by generating complex scenarios whose properties can be finely tuned to challenge and assess autonomous vehicle systems. Experimental results show that our approach has the lowest collision rate of $0.098$ and lowest off-road rate of $0.096$, demonstrating superiority over existing baselines. The proposed approach significantly improves the realism, stability and controllability of the generated scenarios, enabling more nuanced safety evaluation of autonomous vehicles.
@@ III-D Reward-Driven MDP
Figure 2: Generated trajectories in various traffic simulations by CLD. In each scenario, the red line indicates the trajectory of the ego agent, while other colors represent surrounding agents. The generated trajectories in complex intersections and roundabouts show collision-free and traffic rules compliant behaviors, demonstrating the robust and adaptive performance of our approach across diverse driving conditions.
@@ IV-B Results
Figure 2 illustrates the trajectories generated by CLD in various complex scenarios, including challenging intersections and roundabouts. The generated trajectories of the ego agent (as illustrated in red) consistently avoid collisions and strictly adhere to traffic rules, underscoring the robustness and adaptability of our approach under diverse driving conditions.
Table I compares our approach against the baselines under the evaluation metrics outlined in Section IV-A, whereas the baselines’ performance is from [37]. In the experiments, our approach demonstrates impressive performance on both collision avoidance and drivable area following. For the no-collision task, our method achieves a realism score of 0.463 and a failure rate of 0.098—both significantly lower than those of the baselines—indicating that the generated trajectories effectively avoid collisions while preserving realistic driving behavior. Similarly, for the no off-road task, our approach yields a realism score of 0.460 and a failure rate of 0.096, demonstrating that the trajectories consistently remain within the designated drivable area and exhibit strong fidelity to real-world conditions. Together, these results confirm that our framework robustly generates scenarios that meet critical safety and performance criteria.
@@ IV-C Ablation Study
To validate the effectiveness of our training-guided mechanism, we conducted an ablation study comparing the full model (with reward guidance) to a variant without it. Notably, when reward guidance is removed, the realism metric decreases to $0.320$ for both tasks—indicating that the generated trajectories more closely resemble the ground truth. However, this improvement in realism comes at a significant cost: the failure rates increase to $0.446$ and $0.435$, respectively, demonstrating reduced controllability. These results confirm that our approach is essential for achieving a balanced trade-off between realism and controllability.

### v3

@@ Abstract
The validation of autonomous driving systems benefits greatly from the ability to generate scenarios that are both realistic and precisely controllable. Conventional approaches, such as real-world test drives, are not only expensive but also lack the flexibility to capture targeted edge cases for thorough evaluation. To address these challenges, we propose a controllable latent diffusion that guides the training of diffusion models via reinforcement learning to automatically generate a diverse and controllable set of driving scenarios for virtual testing. Our approach removes the reliance on large-scale real-world data by generating complex scenarios whose properties can be finely tuned to challenge and assess autonomous vehicle systems. Experimental results show that our approach has the lowest collision rate of $0.226$ and lowest off-road rate of $0.136$, demonstrating superiority over existing baselines. The proposed approach significantly improves the realism, stability and controllability of the generated scenarios, enabling more nuanced safety evaluation of autonomous vehicles.
@@ III-D Reward-Driven MDP
Figure 2: Generated trajectories in various traffic simulations by CLD. In each scenario, the red line indicates the trajectory of the ego agent, while other colors represent surrounding agents. The generated trajectories in complex intersections show collision-free and traffic rules compliant behaviors, demonstrating the robust and adaptive performance of our approach across diverse driving conditions.
@@ IV-A Experimental Setup
Closed-loop Traffic Simulation. We evaluate our model using closed-loop traffic simulations, where each agent iteratively generates short-horizon trajectory predictions and executes corresponding actions. Simulations are initialized using ground-truth states from the nuScenes dataset and run for 20 seconds with a re-planning frequency of 2 Hz. This closed-loop setup ensures predictions dynamically adapt to evolving environmental conditions, accurately reflecting realistic driving scenarios.
@@ IV-B Results
Figure 2 illustrates the trajectories generated by CLD in various complex scenarios. The generated trajectories of the ego agent (as illustrated in red) consistently avoid collisions and strictly adhere to traffic rules, underscoring the robustness and adaptability of our approach under diverse driving conditions.
Table I compares our approach against the baselines under the evaluation metrics outlined in Section IV-A, whereas the baselines’ performance is from [37]. In the experiments, our approach demonstrates impressive performance on both collision avoidance and drivable area following. For the no-collision task, our method achieves a failure rate of $0.226$ — significantly lower than those of the baselines — indicating that the generated trajectories effectively avoid collisions. Similarly, for the no off-road task, our approach yields a realism score of $0.655$ and a failure rate of $0.136$, demonstrating that the trajectories consistently remain within the designated drivable area and exhibit strong fidelity to real-world conditions. Together, these results confirm that our framework robustly generates scenarios that meet critical safety and performance criteria.
@@ IV-C Ablation Study
To validate the effectiveness of our training-guided mechanism, we conducted an ablation study comparing the full model (with reward guidance) to a variant without it. Notably, when reward guidance is removed, the realism metric decreases to $0.350$ for both tasks — indicating that the generated trajectories more closely resemble the ground truth. However, this improvement in realism comes at a significant cost: the failure rates increase to $0.483$ and $0.450$, respectively, demonstrating reduced controllability. These results confirm that our approach is essential for achieving a balanced trade-off between realism and controllability.